Hydro Systems Engineering Associates, Inc., provides consulting services to city water authorities. The consulting firm’s contribution-margin ratio is 15 percent, and its annual fixed expenses are $285,000. The firm’s income-tax rate is 30 percent.

Required:
1. Calculate the firm’s break-even volume of service revenue.
2. How much before-tax income must the firm earn to make an after-tax net income of $154,000?
3. What level of revenue for consulting services must the firm generate to earn an after-tax net income of $154,000?
4. Suppose the firm’s income-tax rate changes to 20 percent. What will happen to the break-even level of consulting service revenue?
Answer #1

Answers

· [1]

A

Fixed expenses

$285,000

B

CM Ratio

15%

C = A/B

Break even volume of service revenue

$1,900,000

Answer

· [2]

A

After tax net income

$154,000

B

Tax rate

30%

C = A/(100% - 30%)

Before tax income to be earned

$220,000

Answer

· [3]

A

After tax net income

$154,000

B

Hence, before tax income

$220,000

C

Fixed expenses

$285,000

D = B+C

Total contribution margin required

$505,000

E

CM Ratio

15%

F = D/E

Level of revenue required

$3,366,667

Answer

· [4]

The Break even level will STAY THE SAME, even if income tax rate changes to 20%.

A

Fixed expenses

$285,000

B

CM Ratio

15%

C = A/B

Break even volume of service revenue

$1,900,000